Circular No. 9133

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

CBAT Director: Daniel W. E. Green; Room 209; Dept. of Earth and Planetary Sciences; Harvard University; 20 Oxford St.; Cambridge, MA 02138, U.S.A.

cbat@iau.org  cbatiau@eps.harvard.edu  ISSN 0081-0304

                URL http://www.cfa.harvard.edu/iau/cbat.html

COMET C/2010 G1 (BOATTINI) A. Boattini, Lunar and Planetary Laboratory, University of Arizona,

reports his discovery of a comet on Catalina Sky Survey CCD images taken with the 0.68-m Schmidt telescope (discovery observation tabulated below), the object described as diffuse with a strong, condensed coma ∼ 40′′ in size, and a fainter coma out to a diameter ∼ 2′ that fades into the background; the coma is slightly elongated in p.a. 50^{o}–55^{o}. After posting on the Minor Planet Center’s ‘NEOCP’ webpage, other CCD astrometrists have commented on the object’s cometary appearance during Apr. 5.2–6.1 UT, including A. R. Gibbs with the Mt. Lemmon 1.5-m reflector (four co-added 5-s exposures at altitude 20^{o} show a core surrounded by ‘fuzziness’, of size 16′′ × 23′′), R. Ligustri (Udine, Italy, 0.35-m reflector; sixteen 60-s exposures; 30′′ coma), E. Pettarin (Farra d’Isonzo, Italy, 0.40-m reflector; elongated 10′′–15′′ coma), F. Castellani and C. Marangoni (Malcesine, Italy, 0.4-m reflector; 37′′ coma and a 1′16′′ tail in p.a. 70^{o}), P. Birtwhistle (Great Shefford, Berkshire, England, 0.40-m reflector, bright twilight; coma diameter 15′′ and tail 30′′ long in p.a. 65^{o}), L. Buzzi (Varese, Italy, 0.60-m reflector; stacked exposures totally 10 min; condensed coma elongated in p.a. ∼ 35^{o}–40^{o} and extending to a diameter of at least 60′′), and J. E. McGaha (Tucson, AZ, U.S.A., 0.36-m reflector; ten stacked 30-s exposures; 7′′ bright nuclear condensation with a round 20′′ coma and a 10′′-wide tail 45′′ long in p.a. 40^{o}).

2010 UT α_{2000} δ_{2000} Mag. Observer
Apr. 5.12030 4^{h}08^{m}50.^{s}37 +27^{o}39′02′′.3 13.9 Boattini

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2010-G27.

T = 2010 Apr. 6.688 TT  ω = 172.^{o}955
Ω = 287.856 }2000.0
q = 1.20110 AU  i = 76.780

COMET C/2010 F4 (MACHHOLZ) Further to IAUC 9132, images by McGaha (see above) on Mar. 27.5

UT show a bright central nuclear condensation 10′′ in diameter with a round coma 40′′ in diameter. Additional visual total-magnitude and comadiameter estimates: Mar. 27.18, 10.7, 1′.5 (J. J. Gonzalez, Asturias, Spain, 0.20-m reflector); 27.528, 11.3, 3′ (D. Machholz, Colfax, CA, 0.47-m refl.).

2010 April 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green